OpenStudy (superflytyguy):

I really need help! What is the product in simplest form? State any restrictions on the variable. Please show your work. (y^2 / y - 3) x (y^2 - y - 6 / y^2 + y)

OpenStudy (superflytyguy):

\[\frac{ y^2 }{ y -3 } \times \frac{ y^2 -y - 6 }{ y^2 + y }\]

OpenStudy (dtan5457):

I'd factor the y^2-y-6, so you can cross out 2 terms.

OpenStudy (superflytyguy):

y^2 - y = 1 I think So, 1 - 6?

OpenStudy (dtan5457):

Are you referring to the trinomial? It would be (y-3)(y+2)

OpenStudy (superflytyguy):

yes because 3 x 2 = 6 and 3 - 2 = 1

OpenStudy (dtan5457):

-3 times 2=-6 -3+2=-1 y^2-y-6

OpenStudy (superflytyguy):

right

OpenStudy (superflytyguy):

I forgot we are dealing with negatives

OpenStudy (dtan5457):

That's alright Now you have a y-3 and a y-3, which you cross out

OpenStudy (dtan5457):

\[\frac{ y^2(y+2) }{ y^2+y }\]

OpenStudy (dtan5457):

You are now up to there, try to simplify more. (you can )

OpenStudy (superflytyguy):

Okay. \[y^2 \times y = y^3 \]

OpenStudy (superflytyguy):

Now we got y^3 + 2 over y squared plus y

OpenStudy (dtan5457):

Good, now try to simplify the denominator

OpenStudy (dtan5457):

y^2(y+2)=y^3+2y^2

OpenStudy (dtan5457):

(numerator)

OpenStudy (superflytyguy):

Simplify the numerator or the denominator?

OpenStudy (dtan5457):

you already did the numerator. your denominator is y^2+y, try to factor it out.

OpenStudy (superflytyguy):

ok

OpenStudy (superflytyguy):

\[y^2 + y = y (y + y)\]

OpenStudy (dtan5457):

Hmm that's pretty close. but check again. y(y+y)=y^2+y^2

OpenStudy (superflytyguy):

Oh I see now

OpenStudy (dtan5457):

Wanna write out your new simplification?

OpenStudy (dtan5457):

Your almost done

OpenStudy (superflytyguy):

yes, but i'm still trying to simplify y squared plus y. I know it's not that hard but for some reason I can't put my finger on it

OpenStudy (dtan5457):

y^2+y= y(y+1) y times 1=y

OpenStudy (superflytyguy):

Thank you! I didn't think about that

OpenStudy (superflytyguy):

Are there any restrictions on the variable?

OpenStudy (dtan5457):

\[\frac{ y^2(2+y) }{ y(y+1) }\] One more simplification, that y^2 and y can be simplified.

OpenStudy (dtan5457):

\[y^5\div y^3=y^2\] Subtraction of exponents

OpenStudy (superflytyguy):

So that's our answer? y^2?

OpenStudy (dtan5457):

No, that was just an example of the formula, lol Use that to simplify the y^2 on the numerator and y on the botton.

OpenStudy (dtan5457):

*bottom

OpenStudy (superflytyguy):

\[y^2(2 + y) = y(y + 2) + 1 \] right?

OpenStudy (superflytyguy):

No wait that would make it 2y

OpenStudy (dtan5457):

I think your thinking too hard.\[\frac{ y^2(2+y) }{ y(y+1) }=\frac{ y(2+y) }{ (y+1) }\]

OpenStudy (dtan5457):

You can't simplify it any further after that

OpenStudy (superflytyguy):

So this is our restriction of the variable?

OpenStudy (dtan5457):

\[y^2\div y^1=y\]

OpenStudy (superflytyguy):

Sorry I'm good in English not math

OpenStudy (dtan5457):

\[\frac{ y(2+y) }{ (y+1) }\] That answers into the most simplified product of your original question

OpenStudy (dtan5457):

As for restrictions, that's asking what numbers can't be used

OpenStudy (dtan5457):

As for as I'm concerned the only restriction for y would be -1, because the denominator is "illegal" if it's 0

OpenStudy (superflytyguy):

Okay I see

OpenStudy (dtan5457):

So I believe that answers all your questions?

OpenStudy (superflytyguy):

I think so! Thank you so much!

OpenStudy (dtan5457):

No problem. Feel free to fan and tag me into more questions. Thanks for cooperating through my "lesson"